How much 0.1200 M sodium hydroxide solution is need to titrate 13 mL of a 0.120 M sulfuric acid solution? Answer with 2 decimal places.
1
Expert's answer
2019-10-02T08:01:16-0400

n(H2SO4) = C*V= 0.12*13

n(NaOH) = 2n (H2SO4)= 2*0.12*13

V(NaOH)=n/C=2*0.12*13/0.12=26.00 mL

The answer is 26.00 mL